find the solution set of y=7x-12 and 4x-2y=34 using elimination, substitution and graphical method.
Answer by satyareddy22(84)

y=7x-12------(1)
4x-2y=34 ----(2)
substituting (1) in (2)
4x-2(7x-12)=34
4x-14x+24=34
-10x=34-24=10
-10x=10
x=10/(-10)=-1=>x=-1
now substituting x=-1 in (1)
y=7(-1)-12=-7-12=-19
y=-19